Please help me with this one...
Number problems. Find two consecutive positive integers such that the sum of their square is 85.
Thank you!

Answer by checkley71(8403)

X^2+(X+1)^2=85
X^2+X^2+2X+1=85
2X^2+2X+1-85=0
2X^2+2X-84=0
2(X^2+X-42)+0
2(X+7)(X-6)=0
X+7=0
X=-7 ANSWER.
X+1=-6 ANSWER.
-------------------------
X=6 ANSWER.
X+1=7 ANSWER.
PROOF
(-7)^2+(-6)^2=85
36+49=85
85=85
PROOF
6^2=7^2=85
36+49=85
85=85